OpenStudy (anonymous):

Medals!! What is the solution to the system of equations? 2x+y=-3 x-2y=-4

OpenStudy (alexandervonhumboldt2):

y=-3-2x y=x/2+2

OpenStudy (alexandervonhumboldt2):

1/2x+2=-3-2x

OpenStudy (alexandervonhumboldt2):

5/2x=-5 x=-2

OpenStudy (alexandervonhumboldt2):

so x=-2 y=1

OpenStudy (alexandervonhumboldt2):

(-2, 1)

OpenStudy (michele_laino):

please, add 2y to both sides of your second equation